OpenStudy (darkigloo):

If y=tanu, u=v-(1/v), and v=lnx, what is dy/dx at x=e?

OpenStudy (darkigloo):

So it will look like \[y = \tan ( lnx - \frac{ 1 }{ lnx })\]

OpenStudy (ipwnbunnies):

Holy crap man. Well, he's the plan. Plug v into u. Plug u into y. Derivative of tan(u) = sec^2(u)*du/dx

OpenStudy (anonymous):

dy/dx = dy/du * du/dv *dv/dx

OpenStudy (ipwnbunnies):

^^ Man, that's great.

OpenStudy (anonymous):

find dy/du, du/dv, and dv/dx and then multiply the three together